# Lightwave-controlled relativistic plasma mirrors

### v2

@@ Abstract
We report on attosecond-scale control of high-harmonic and electron emission from plasma mirrors driven by relativistic-intensity near-single-cycle lightwaves at kHz repetition rate. By controlling the waveform of the intense light transient, we reproducibly form a sub-cycle temporal intensity gate at the plasma mirror surface, leading to the observation of extreme ultraviolet spectral continua, characteristic of isolated attosecond pulse generation. We also observe the correlated emission of a waveform-dependent relativistic electron beam, paving the way towards fully lightwave-controlled dynamics of relativistic plasma mirrors.
The generation of attosecond light pulses has enabled the real-time observation of the fastest electronic processes in matter. Although new generation schemes have emerged [1, 2], the currently preferred route to produce attosecond pulses remains High-Harmonic Generation (HHG) from the gas phase [3]. The mechanism is now well understood, the technology is mature (as indicated by the emergence of turn-key, tabletop HHG beamlines), and it has been successfully used in several pioneering applications unveiling previously unobservable attosecond-scale dynamics [4, 5, 6]. However, it is intrinsically limited to low intensity ($\sim 10^{15}\,\rm W/cm^{2}$) driving pulses and the maximum laser-to-XUV conversion efficiency is currently of the order of $10^{-4}$ [7, 8]. To circumvent those limitations, HHG from relativistic plasma mirrors has long been recognized as a promising route [9] owing to the predicted percent-level conversion efficiencies [10] (although the measured efficiencies are $\sim 10^{-4}$ so far [11, 12]), their compatibility with relativistic-intensity ($>10^{18}\,\rm W/cm^{2}$) driving lasers, as well as the production of chirp-free attosecond pulses [13] with high spatial coherence [14, 15].
@@
Experiments were conducted using the Salle Noire laser driver [31], developed in-house at Laboratoire d’Optique Appliquée, which routinely delivers multi-mJ energy sub-4 fs pulses (FWHM) at a central wavelength $\lambda_{0}\approx 780$ nm, corresponding to less than 1.5 optical cycles, with a temporal contrast ratio $>10^{10}$ up to 15 ps and better than $10^{7}$ up to 3 ps before the pulse peak. A shot-to-shot CEP stability of $\approx$ 380 mrad rms can be maintained for extended ($\sim$ hour) periods of time [32]. The CEP, which is always inherently measured relative to an unknown offset value, can thus be varied at will while keeping all other laser parameters identical.
@@
Figure 2: Vertically-integrated HHG spectra as a function of relative CEP for Prepulse lead times of: -1 ps (a), 1 ps (b), and 2 ps (c). Above each frame are two line-out spectra measured for opposite laser waveforms ($\pi$ CEP offset). Each data acquisition is averaged over 100 consecutive laser shots (100 ms). To the right of each frame, we show the temporal auto-correlation of the HHG emission in the shown spectral range. The white dashed line marks the position of the satellite peak near 1 laser period. The scans were made in succession while the CEP-locking loop remained active. The relative CEPs thus correspond to the same driving waveforms in each panel.
@@
Figure 3: Simultaneous acquisition of XUV spectra (top panel), electron beam profiles (middle panel), and electron spectra (bottom panel) for different relative CEP values. Each data acquisition is averaged over 100 consecutive laser shots (100 ms). The red dot in the middle panel indicates the point of measurement of the electron spectrum. The dashed boxes indicate the two opposite waveforms in the electron push-pull scenario: at the top, the angle-integrated HHG spectrum is shown for the two opposite driving waveforms and corresponding push-pull scenarios at absolute CEP values, $\psi_{\textrm{CEP}}=\pi/2$ and $23\pi/$ 2. The formation of an XUV continuum can clearly be resolved in the case of a single push-pull sequence for a given relative CEP (step size $<0.5\>$rad).
@@
In order to estimate the surface plasma gradient scale length $L_{g}=L_{0}+c_{s}\tau$, where $L_{0}\approx 0.01\lambda_{0}$ is the estimated contribution of the Main-pulse finite temporal contrast, and $\tau$ the delay between the Prepulse and the Main-pulse, we experimentally measured the plasma expansion speed $c_{s}$ using spatial domain interferometry [35, 33] and found $c_{s}=22$nm/ps. In fig. 2, we show the vertically-integrated XUV spectra recorded as a function of relative CEP for three different values of $\tau$: -1 ps, 1 ps, and 2 ps, corresponding to respectively $L_{g}\approx L_{0}$, $0.04\lambda_{0}$ and $0.07\lambda_{0}$. We observe clear CEP and $L_{g}$-dependent spectral structures of the harmonics. Assuming a nearly flat spectral phase [13], these can be linked to the periodicity of the corresponding attosecond temporal profiles, apparent in the temporal auto-correlations obtained as the modulus of the Fourier transform of each HHG spectrum in the shown spectral range and also shown in fig. 2. For $L_{g}\approx L_{0}$ (fig. 2a), the slower spectral modulation on top of the harmonic orders probably results from the presence of both coherent wake emission [36] and relativistic HHG. Their different emission times [13] then create 2 attosecond pulses per $\approx 2.5$-fs laser period, which appears as an additional peak near 1.3 fs in the auto-correlations. As the plasma gradient softens to $L_{g}\approx 0.04\lambda_{0}$ (fig. 2b), relativistic HHG becomes dominant, and only the satellite peak near the $\approx 2.5$-fs laser period remains in the auto-correlation. Its CEP-dependent shift over $\approx 0.3\>$fs likely results from a CEP-dependent plasma denting evolution over the pulse duration [12]. Finally, in panel (c), there’s interestingly no clear temporal shift of the satellite peak (i.e., the harmonic modulation period is nearly constant).
@@
Finally, we find that a 2-ps prepulse delay ($L_{g}\approx 0.07\lambda_{0}$, fig. 2c) maximizes the spectral extent of the harmonics (from $\tau=3\>$ps on, both harmonic signal yield and spectral extension drop), which is close to the gradient optimum of $0.1\lambda_{0}$ typically observed for relativistic HHG [37]. Only for this optimum plasma gradient scale length do we observe clear XUV spectral continua for a given relative CEP value (000\>rad in fig.2c), which is characteristic of IAP generation. Clearly modulated spectra are observed when the relative CEP is offset by $\pi$. Interestingly, the corresponding $\approx 2.5$-fs satellite peak in the temporal auto-correlations shifts by no more than $\approx 0.1\>$fs before falling below the noise level.
@@
Figure 4: Angular electron charge distribution scanned over randomized CEP values, measured at the $\tau=2\>$ps delay between the Main-pulse and Prepulse. Each electron image is averaged over 100 consecutive laser shots (100 ms). The dashed black line indicates the specular direction at $\phi_{az}=55^{\circ}$ w.r.t. target normal.
@@
Measuring the electron energy spectrum (point of measurement indicated by the red dot in the middle panel of fig. 3) confirms the clear influence of the CEP on the electron acceleration mechanism. The dashed boxes in fig. 3 highlight the acquired electron spectra under similar conditions for the two opposite laser waveforms: In the right box, for a given relative CEP (000\>rad), we are in the scenario in which the push-pull mechanism only occurs once, as indicated by the production of an XUV continuum. Interestingly, this corresponds to the production of an electron energy spectrum with a single broad peak at $\approx$ 1 MeV. As the CEP is shifted by $\pi$ (left box, CEP$=-3~{}$rad), a more modulated electron spectrum is observed, with the appearance of a shoulder at lower energies. These features could be the indication of the phase-locked injection and subsequent VLA of either one (in the optimal-gating case of $\psi_{\textrm{CEP}}=3\pi/2$) or two (in the case of $\psi_{\textrm{CEP}}=\pi/2$) attosecond relativistic electron bunches. These preliminary results should be confirmed in a more detailed experimental study. Unfortunately, our setup only allowed us to measure the spectrum at the periphery of the electron beam, as indicated by the red dot in the middle panel in fig. 3. We expect higher energies in the beam center, in particular in the optimal-gating scenario (right box), and intend to confirm this in future experiments with a new design for the diagnostics. Our observation is nonetheless the first experimental evidence of CEP-controlled emission of relativistic electrons from plasma mirrors.
@@
Fig. 4 shows the full electron beam measured under similar conditions with the wide-angle Lanex screen, as the relative CEP is cycled through $2\pi$. The electron angular charge distribution exhibits the same CEP dependence as in the previous measurements (middle panel in fig. 3), with maximum charge peak appearing at a relative CEP of 000~{}rad and a minimally charged beam (approximately half the total integrated charge) when the CEP is shifted by $\pi$. It is interesting to note that although the distribution and direction of the electron beam do not change significantly with CEP, there is, however, a small sign of ponderomotive blowout around the specular direction in the case of relative CEP of 3 rad, which would indicate that when $\psi_{\textrm{CEP}}=\pi/2$ one of the electron bunches is not injected with the adequate phase-energy combination for efficient VLA. We plan to measure the spatially-resolved electron spectrum across the full beam in order to visualize CEP-dependent fine structure that could provide more insight into the attosecond-scale electron acceleration dynamics.
@@ Acknowledgements.
This work was supported by the Agence Nationale pour la Recherche (ANR-11-EQPX-005-ATTOLAB, ANR-14-CE32-0011-03 APERO, ANR-22-CE30-0005-01 BANDITO); European Research Council (ERC FEMTOELEC 306708, ERC ExCoMet 694596); LASERLAB-EUROPE (H2020-EU.1.4.1.2. grant agreement ID 654148), Région Ile-de-France (SESAME 2012-ATTOLITE).

### v3

@@ Abstract
We report on attosecond-scale control of high-harmonic and fast electron emission from plasma mirrors driven by relativistic-intensity near-single-cycle lightwaves at kHz repetition rate. By controlling the waveform of the intense light transient, we reproducibly form a sub-cycle temporal intensity gate at the plasma mirror surface, leading to the observation of extreme ultraviolet spectral continua, characteristic of isolated attosecond pulse generation. We also observe the correlated emission of a waveform-dependent relativistic electron beam, paving the way from towards fully lightwave-controlled dynamics of relativistic plasma mirrors.
The generation of attosecond light pulses has enabled the real-time observation of the fastest electronic processes in matter. Although new generation schemes have emerged [1, 2], the currently preferred route to produce attosecond pulses remains High-Harmonic Generation (HHG) from the gas phase [3, 4]. The mechanism is now well understood, the technology is mature (as indicated by the emergence of turn-key, tabletop HHG beamlines), and it has been successfully used in several pioneering applications unveiling previously unobservable attosecond-scale dynamics [5]. However, it is intrinsically limited to low intensity ($\sim 10^{15}\,\rm W/cm^{2}$) driving pulses and the maximum laser-to-XUV conversion efficiency is currently $\lesssim 10^{-4}$ [3, 4]. To circumvent those limitations, HHG from relativistic plasma mirrors has long been recognized as a promising route [6] owing to the predicted percent-level conversion efficiencies [7] (although the measured efficiencies are $\sim 10^{-4}$ so far [8, 9]), their compatibility with relativistic-intensity ($>10^{18}\,\rm W/cm^{2}$) driving lasers, as well as the production of chirp-free attosecond pulses [10] with high spatial coherence [11, 12].
@@
Experiments were conducted using the Salle Noire laser driver [28], developed in-house at LOA , which routinely delivers multi-mJ energy sub-4 fs pulses (FWHM) at a central wavelength $\lambda_{0}\approx 780$ nm, corresponding to less than 1.5 optical cycles, with a temporal contrast ratio $>10^{10}$ up to 15 ps and better than $10^{7}$ up to 3 ps before the pulse peak. A shot-to-shot CEP stability of $\approx$ 380 mrad rms can be maintained for $\sim$ hour durations [29]. The CEP is measured and locked relative to an unknown offset value at the 1-kHz laser repetition rate by a Fringeezz device with feedback to a Dazzler (both Fastlite). It can be varied while keeping all other laser parameters identical.
@@
Figure 2: Vertically-integrated HHG spectra as a function of relative CEP for Prepulse lead times of: -1 ps (a), 1 ps (b), and 2 ps (c). Above each frame are two line-out spectra measured for opposite laser waveforms ($\pi$ CEP offset). Each acquisition averaged over 100 consecutive laser shots (100 ms). To the right of each frame, we show the temporal auto-correlation of the HHG emission in the shown spectral range. The white dashed line marks the position of the satellite peak near 1 laser period where such a peak could be determined. The scans were made in succession while the CEP-locking loop remained active. Equal relative CEPs thus correspond to the same driving waveforms in each panel.
@@
Figure 3: Simultaneous acquisition of XUV spectra (top panel), electron beam profiles (middle panel), and electron spectra (bottom panel) for different relative CEP values. Each acquisition averaged over 100 consecutive laser shots (100 ms). The red dot in the middle panel indicates the point of measurement of the electron spectrum. The dashed boxes indicate the two opposite waveforms in the electron push-pull scenario: at the top, the angle-integrated HHG spectrum is shown for the two opposite driving waveforms and corresponding push-pull scenarios at absolute CEP values, $\psi_{\textrm{CEP}}=\pi/2$ and $23\pi/$ 2. The formation of an XUV continuum can clearly be resolved in the case of a single push-pull sequence for a given relative CEP (step size $<0.5\>$rad).
@@
In order to estimate the surface plasma gradient scale length $L_{g}=L_{0}+c_{s}\tau$, where $L_{0}\approx 0.01\lambda_{0}$ is the estimated contribution of the Main-pulse finite temporal contrast, and $\tau$ the delay between the Prepulse and the Main-pulse, we experimentally measured the plasma expansion speed $c_{s}$ using spatial domain interferometry [32, 30] and found $c_{s}=22$nm/ps. In fig. 2, we show the vertically-integrated XUV spectra recorded as a function of relative CEP for three different values of $\tau$: -1 ps, 1 ps, and 2 ps, corresponding to respectively $L_{g}\approx L_{0}$, $0.04\lambda_{0}$ and $0.07\lambda_{0}$. We observe clear CEP and $L_{g}$-dependent spectral structures of the harmonics. Assuming a nearly flat spectral phase [10], these can be linked to the periodicity of the corresponding attosecond temporal profiles, apparent in the temporal auto-correlations obtained as the modulus of the Fourier transform of each HHG spectrum in the shown spectral range and also shown in fig. 2. For $L_{g}\approx L_{0}$ (fig. 2a), the slower spectral modulation on top of the harmonic orders probably results from the presence of both coherent wake emission [33] and relativistic HHG. Their different emission times [10] then create 2 attosecond pulses per $\approx 2.5$-fs laser period, which appears as an additional peak near 1.3 fs in the auto-correlations. As the plasma gradient softens to $L_{g}\approx 0.04\lambda_{0}$ (fig. 2b), relativistic HHG becomes dominant, and only the satellite peak near the $\approx 2.5$-fs laser period remains in the auto-correlation. Its CEP-dependent shift over $\approx 0.3\>$fs likely results from a CEP-dependent plasma denting evolution over the pulse duration [9]. Finally, in panel (c), there is interestingly no clear temporal shift of the satellite peak (i.e., the harmonic modulation period is nearly constant).
@@
Finally, we find that a 2-ps prepulse delay ($L_{g}\approx 0.07\lambda_{0}$, Fig. 2c) maximizes the spectral extent (for $\tau>3\>$ps, both harmonic yield and spectral extent drop), which is close to the optimal gradient of $L_{g}\approx 0.1\lambda_{0}$ typically observed for relativistic HHG [34, 24]. Only for this optimum gradient do we observe clear XUV spectral continua for a given relative CEP (000\>rad in Fig.2c), characteristic of IAP generation. Clearly modulated spectra are observed when the relative CEP is offset by $\pi$. Interestingly, the corresponding $\approx 2.5$-fs satellite peak in the temporal auto-correlations shifts by no more than $\approx 0.1\>$fs before falling below the noise level.
@@
Figure 4: Angular electron charge distribution for the locked CEP values indicated in the panels, measured at $\tau=2\>$ps Prepulse lead time. Each image averaged over 100 consecutive laser shots (100 ms). Red ticks indicate the specular direction.
@@
Measuring the electron energy spectrum (point of measurement indicated by the red dot in the middle panel of fig. 3) confirms the clear influence of the CEP on the electron acceleration mechanism. The dashed boxes in fig. 3 highlight the acquired electron spectra under similar conditions for the two opposite laser waveforms: In the right box, for a given relative CEP (000\>rad), we are in the scenario in which the push-pull mechanism only occurs once, as indicated by the production of an XUV continuum. Interestingly, this corresponds to the production of an electron energy spectrum with a single broad peak at $\approx$ 1 MeV. As the CEP is shifted by $\pi$ (left box, CEP$=-3~{}$rad), a more modulated electron spectrum is observed, with the appearance of a shoulder at lower energies. These features could be the indication of the phase-locked injection and subsequent VLA of either one (in the optimal-gating case of $\psi_{\textrm{CEP}}=3\pi/2$) or two (in the case of $\psi_{\textrm{CEP}}=\pi/2$) attosecond relativistic electron bunches. In the latter case, one bunch undergoes less efficient VLA than the other. These preliminary results should be confirmed in a more detailed experimental study. Unfortunately, our setup only allowed us to measure the spectrum at the electron beam periphery, indicated by the red dot in the middle panel in Fig. 3. We expect higher energies in the beam center, in particular in the optimal-gating scenario (right box). Our observation is nonetheless the first experimental evidence of CEP-controlled emission of relativistic electrons from plasma mirrors.
@@
Fig. 4 shows the full electron beam measured under similar conditions with the wide-angle Lanex screen, as the relative CEP is cycled through $2\pi$. The electron angular charge distribution exhibits the same CEP dependence as in the previous measurements (middle panel in fig. 3), with maximum charge peak appearing at a relative CEP of 000~{}rad and a minimally charged beam (approximately half the total integrated charge) when the CEP is shifted by $\pi$. Although the distribution and direction of the electron beam only subtly changes, for a relative CEP of 000~{}rad, the less divergent “VLA beam”, centered on $\phi_{\mathrm{az}}\approx 50^{\circ}$, is most dominant over the more divergent contribution centered near the specular direction. This dominance gradually fades with CEP until for a relative CEP of 3 rad, a small sign of ponderomotive blowout appears around the specular direction. This would indicate that when $\psi_{\textrm{CEP}}=\pi/2$, one of the electron bunches is not injected with the adequate phase-energy combination for efficient VLA, leading to a lower final energy and some ponderomotive scattering. We plan to measure the spatially-resolved electron spectrum across the full beam in order to visualize CEP-dependent fine structure that could provide more insight into the attosecond-scale electron acceleration dynamics.
@@
Data underlying the results presented in this paper are available in Ref. [35].
@@ Acknowledgements.
This work was supported by the Agence Nationale pour la Recherche (ANR-11-EQPX-005 ATTOLAB, ANR-14-CE32-0011-03 APERO, ANR-22-CE30-0005-01 BANDITO); European Research Council (ERC FEMTOELEC 306708, ERC ExCoMet 694596); LASERLAB-EUROPE (H2020-EU.1.4.1.2. grant agreement ID 654148), Région Ile-de-France (SESAME 2012-ATTOLITE).